OpenStudy (anonymous):

The circumference of a circle is 28x(pi) meters.what is its radius? _____i know I'm supposed to use 3.14. But my question is do I multiply 3.14x28?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

and no

OpenStudy (misty1212):

the circumference is \(2\pi r=28\pi\) divide by \(2\pi\) to find the radius, i.e. find \[\frac{28\pi}{2\pi}\]

OpenStudy (anonymous):

So 14?